OpenStudy (anonymous):

Write two linear equations you can use to graph y=|x+7|

OpenStudy (anonymous):

When x >= -7, then x + 7 >= 0 so you would have for part of the solution: y = x + 7 for x >= -7 When x < -7, then x + 7 < 0 so the other part of the solution is: y = -(x + 7) = -x - 7 for x < -7 That covers all values of "x", so depending on where you are horizontally, you use that particular equation.

OpenStudy (anonymous):

All good now @Kagome_gurl8 ?

OpenStudy (anonymous):

Yes thank you so much :D

OpenStudy (anonymous):

Hold on a minute, I've got something helpful.

OpenStudy (anonymous):

Here's a graph for you.

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Kagome_gurl8 And you're very welcome. Nice working with you.

OpenStudy (anonymous):

So it'd be